# Perfect Wave Transfer in Continuous Quantum Systems

### v1

@@ Abstract
We study the perfect transfer of information in 1+1D continuous quantum systems. This includes effective descriptions of inhomogeneous spin chains, for which the notion of perfect state transfer in quantum information was introduced, and here phrased in terms of waves. We show that reflection symmetry is necessary for perfect wave transfer (PWT) in any inhomogeneous conformal field theory, and even sufficient when restricted to one-particle excitations. To determine if or when it is sufficient more generally, we first break conformal invariance and study a broad class of 1+1D bosonic theories. We show that the question can then be posed as an inverse Sturm-Liouville problem that determines when the bosonic theory exhibits PWT. We demonstrate how to uniquely solve this problem, which also shows that reflection symmetry is sufficient for the special case with conformal invariance. Using bosonization, our continuum results extend these notions to interacting quantum field theories.
Introduction.—Non-photonic quantum communication enables the efficient transmission of information from one quantum register to another [1]. This is desirable for quantum networks and computation [2, 3], not least since it allows for the perfect transfer of states even without dynamical control [4, 5]. First shown theoretically for communication channels in the form of a spin chain with spatially varying hopping, such perfect state transfer (PST) was extended to spins on graphs and other networks [6, 7, 8, 9, 10, 11, 12, 13, 14] and demonstrated experimentally [15, 16, 17, 18, 19, 20]. However, the notion has remained squarely in the discrete realm.
@@
In this paper, we extend PST to continuous systems in one space and one time dimension (1+1D). In so doing, a rephrasing in terms of waves becomes natural, motivating the term perfect wave transfer (PWT). As for states in PST, a system on the interval $[-L/2,L/2]$ of length $L$ with suitable boundary conditions (BCs) exhibits PWT if the propagation of any initial wave, after some time $T$, amounts to its reflection; see Fig. 1. Concretely: The system enjoys PWT if
$\langle\mathcal{O}(x,0)\rangle=\langle\mathcal{O}(-x,T)\rangle,$ (1)
@@
Related recent advances in the continuum include non-equilibrium studies of inhomogeneous versions of conformal field theory (CFT). These are conformally-invariant quantum field theories deformed so that the usual constant propagation velocity becomes a profile $v(x)$ [21, 22, 23, 24, 25, 26]. Among their applications, such inhomogeneous CFTs effectively describe gapless quantum many-body systems with mesoscopic inhomogeneities, such as spin chains with spatially varying hopping, trapped ultracold atoms, and arrays of quantum circuits; cf. [27]. Special cases include the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$ for some constant $v>0$ [28, 29]—the continuum counterpart of the inhomogeneous spin chain in [4] that first exhibited PST—and profiles produced by Möbius transformations [30, 31, 32, 33, 34, 35, 36, 37]. Note that for PWT, as for other transport properties of finite systems, the BCs [38, 39, 29, 40] are crucial. Furthermore, inhomogeneous CFT can be viewed as a theory in curved spacetime. As such, besides applications to quantum communication in solid-state devices [41, 42], our work also relates to relativistic quantum information [43, 44]. Since the same theories are also natural to study on the half or full infinite line, our results may likewise be relevant for long-distance quantum communication [45].
Figure 1: (a) Illustration of PWT with Neumann BCs showing an initial wave that evolves to its reflection at a certain time $T$. (b) Its manifestation in an inhomogeneous CFT with a given profile $v(x)/v_{0}$ (inset), showing a localized initial wave propagating along curved light-cone trajectories that recombine at the reflected point at $t=T\equiv L/v_{0}$ given by Eq. (7). Specifically, the magnitude of boson two-point correlations (17) at the free-fermion point is plotted, which by bosonization encodes the same information as $\log F(t)$ in Eq. (5) for delta-like waves $\xi_{1}$ centered at $-3L/8$ and $\xi_{2}$ at $x$ (setting $k_{\textrm{F}}=0$ for simplicity and $t^{\prime}=0$).
Engineering PST/PWT is at heart a spectral problem, amounting to whether the propagator coincides with the parity operator at certain times. We address this engineering question for quantum field theories describing a continuous quantum communication channel. Remarkably, for bosonic theories [46, 47, 48, 49, 50, 51, 52, 53], we arrive at an inverse problem for Sturm-Liouville operators, which we solve and whose solution determines if the channel enjoys PWT. Using bosonization [54, 55, 56, 57, 58, 59], our results extend the notion of PST/PWT to a large class of systems, including interacting fermionic quantum field theories.
Inhomogeneous CFT.—To begin with, we study the continuum limit of the inhomogeneous spin chains in [1, 4]. In the gapless regime, the continuous counterpart is an inhomogeneous CFT with a velocity profile $v(x)>0$, $x\in[-L/2,L/2]$ [21, 22, 23, 24, 25, 26]. The simplest example is that of inhomogeneous free fermions, given by the Hamiltonian (up to an overall constant and setting $\hbar=1$)
$H_{\textrm{iFF}}\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\frac{v(x)}{2}\Bigl{[}\!\left.:\!\psi^{\dagger}_{+}(-\mathrm{i})\partial_{x}\psi^{\vphantom{\dagger}}_{+}\!:\right.+\left.:\!\psi^{\dagger}_{-}\mathrm{i}\partial_{x}\psi^{\vphantom{\dagger}}_{-}\!:\right.+\mathrm{h.c.}\Bigr{]}$ (2)
[…]
We assume open (Neumann) BCs at $x=\pm L/2$. These can be incorporated by standard unfolding methods; cf. [61, 62, 57, 58, 38, 63]: Couple right and left movers by identifying $\psi_{-}(x)=\psi_{+}(L-x)$, giving a chiral theory of only right-moving fermions on $[-L/2,3L/2]$ with periodic BCs and velocity $v(x)$ for $x\in[-L/2,L/2]$ and $v(L-x)$ for $x\in[L/2,3L/2]$. This implies zero particle flow at $x=\pm L/2$ since the two fermion densities $\rho_{\pm}(x)=\left.:\!\psi^{\dagger}_{\pm}(x)\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.$ are equal at each boundary and the current is $\propto\rho_{+}-\rho_{-}$.
PST was probed in [4] using a time-dependent overlap between one-particle states that are reflections of each other. Its continuum counterpart is to study states created by fields
@@
where $k_{\textrm{F}}$ denotes the Fermi momentum. Eq. (3) corresponds to the low-energy linearization of the spin-chain dispersion relation around its Fermi points ±$\pm k_{\textrm{F}}$, followed by recentering the resulting fields so they describe right/left-moving excitations around zero momentum. As a preparatory subset of possible states, consider
@@
for smooth enough functions $\xi_{j}(x)$, interpreted as distributions or “wave packets” of fermionic excitations in each state, and study the time-dependent overlap
@@
The states (4) form a restricted class as only one-particle excitations are possible and as right/left movers have the same particle distributions; cf. the more general treatment in the SM [60] where we lift the latter restriction.
To probe PWT for one-particle excitations as done in [4], we restrict to specular waves, $\xi_{2}(x)=\xi_{1}(-x)$, and compare $F(t)$ in Eq. (5) with $\langle\Psi_{1}|\Psi_{1}\rangle$. We say that: There is one-particle PWT if there exists $T>0$ so that
@@
for all initial waves $\xi_{1}(x)$ in Eq. (4)  [^1] If Eq. (6) holds for any state $|\Psi_{1}\rangle$ in the Fock space, then general PST/PWT (i.e., for any number of excitations) would be enjoyed..
Eq. (6) amounts to a comparison between two-point correlation functions at equal and non-equal times. Such can be computed in any inhomogeneous CFT by mapping them to the corresponding homogeneous correlations [26]. The details are given in the SM [60] and boil down to comparing spacetime points after conformal transformations to coordinates given by
@@
The outcome of the comparison shows that reflection symmetry is essential:
An inhomogeneous CFT exhibits one-particle PWT
if and only if $v(x)$ is even.
The corresponding times are odd multiples, $(2m+1)T$ for any integer $m$, of the time $T=L/v_{0}$ that excitations take to cross the entire system’s length. For later usage, we note that $y\in[-L/2,L/2]$ if $v(x)$ is even.
@@
As an example, consider the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$, which gave the effective low-energy theory of the spin chain in [4]. This continuum theory exhibits PWT, just like for any even profile $v(x)$, in contrast to [4] where that profile was exceptional in this respect. Notably, the square-root profile is positive everywhere except at $x=\pm L/2$. Thus, strictly speaking, $v(x)$ should be regularized before using the tools in [26], but since the integral defining $1/v_{0}$ in Eq. (7) converges, it has no consequence for PWT. Oppositely, the limiting case of inhomogeneous CFT given by Möbius transformations so that $v(x)$ decays to zero at the boundaries cannot exhibit PWT since $1/v_{0}$ diverges.
@@
Bosonic theories.—The above result and observations should not come as a surprise. After all, if one tries to compare discrete and continuous theories, then an inhomogeneous CFT can only describe low-energy excitations of a given lattice model. Moreover, one can show [60] that any Hamiltonian (with pure point spectrum) must commute with the parity operator for PWT to be possible, i.e., $v(x)$ must be even. The general structure of the spectrum due to conformal invariance in 1+1D [65] opens up the possibility that it might also be sufficient; e.g., the example (2) has a linear spectrum and is one-particle reducible, suggesting that if PWT holds for one-particle excitations then this should extend to many particles. The remainder of this paper concerns to what extent this is true more generally. Namely: What are necessary and sufficient conditions for PWT (1) for any number of excitations?
@@
To answer this question, we first break conformal invariance. Examples include 1+1D fermions with non-local interactions [66, 67, 68], while here we focus on bosons, specifically inhomogeneous Tomonaga-Luttinger liquids (TLLs) [46, 47, 48, 49, 50, 51, 52, 53]. This theory is formulated using a compactified bosonic field $\varphi(x)$ (modulo $2\pi$) with conjugate $\Pi(x)$ satisfying $[\partial_{x}\varphi(x),\Pi(x^{\prime})]=\mathrm{i}\partial_{x}\delta(x-x^{\prime})$. The Hamiltonian is (up to an overall constant)
@@
$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}\biggr{)}\!:\right.\!,$ (8)
@@
where, besides $v(x)>0$, the so-called Luttinger parameter $K(x)>0$ also depends on $x$. For simplicity, we assume that both functions are smooth. When $K(x)=K$ is constant, Eq. (8) is an inhomogeneous CFT, and when $K=1$, it is the bosonized version of the free fermions (2) [54, 55, 56, 57, 58, 59]. However, $K^{\prime}(x)\neq 0$ locally couples right and left movers non-trivially, thus they do not commute and the theory is not conformal [53]. Again, we impose open BCs, namely zero current at the boundaries:
@@
$\jmath(x)\big{|}_{x=\pm L/2}=0,\quad\textnormal{where}\;\;\jmath(x)\equiv-\frac{v(x)K(x)}{\pi}\partial_{x}\varphi(x)$ (9)
@@
is the total particle current. Unless stated otherwise, we also assume that $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$ so that right and left movers are only coupled via the BCs (9) at those points.
@@
Sturm-Liouville (SL) theory.—Following [69, 70, 71, 72, 73, 74, 27], we use SL theory to find eigenfunctions, using which the fields can be expanded and the Hamiltonian (8) diagonalized. (This generalizes [27] as $v(x),K(x)$ are unrelated and differs from the approach in [53].) Indeed, using integration by parts and Eq. (9),
@@
$H_{\textrm{iTLL}}=\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}w(x)\!\left.:\!\Biggl{(}\biggl{[}\frac{\pi\Pi(x)}{w(x)}\biggr{]}^{2}+\varphi(x)\mathcal{A}\varphi(x)\Biggr{)}\!:\right.\!$ (10)
@@
$w(x)\equiv\frac{K(x)}{v(x)},\quad p(x)\equiv v(x)K(x),\quad q(x)\equiv 0,$ (12)
where $\mathcal{A}$ acts on the space $\mathcal{D}$ of smooth  [^2] More precisely, twice continuously differentiable. functions $u(x)$ on $[-L/2,L/2]$ that satisfy $v(x)K(x)\partial_{x}u(x)\big{|}_{x=\pm L/2}=0$, i.e., Neumann BCs when $v(x),K(x)>0$.
@@
A distinction is warranted: Eq. (13) is regular if certain technical conditions are satisfied  [^3] Here, these amount to $v(x)$, $K(x)$ being strictly positive and $v(x)$, $K(x)$, $\partial_{x}p(x)$ [and $q(x)$ if non-zero] continuous on the whole interval $[-L/2,L/2]$.; otherwise it is irregular. Its importance stems from properties that hold for regular problems [77]:
@@
    The spectrum consists of eigenvalues $\lambda_{n}$ that are real and ordered: $0\leq\lambda_{0}<\lambda_{1}<\ldots$.
@@
    Each $\lambda_{n}$ has a unique square-integrable eigenfunction $u_{n}(x)$ with $n$ isolated zeros in $[-L/2,L/2]$.
@@
    The $u_{n}(x)$ form a complete orthonormal basis for square-integrable functions on $[-L/2,L/2]$ with weight $w(x)$: $\int_{-L/2}^{L/2}\mathrm{d}x\,w(x)u_{n}(x)u_{n^{\prime}}(x)=\delta_{n,n^{\prime}}$.
Furthermore, assuming that the integral defining $1/v_{0}$ in Eq. (7) converges:
@@
    The $\lambda_{n}$ can at most grow as $n^{2}$ for large $n$, i.e., $\lambda_{n}=O(n^{2})$.
The last follows by Weyl’s law; see the SM [60] and, e.g., [78, 79]. Simply put, it gives a semi-classical estimate of the number of eigenvalues below a given large $\lambda>0$ for Laplace-type operators.
@@
$\varphi(x)=\sum_{n=0}^{\infty}\varphi_{n}u_{n}(x),\qquad\Pi(x)=\sum_{n=0}^{\infty}\Pi_{n}w(x)u_{n}(x),$ (14)
whereby we obtain $\bigl{[}\varphi_{n^{\vphantom{\prime}}},\Pi_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$ and can diagonalize $H_{\textrm{iTLL}}$ in Eq. (8). Indeed, the latter becomes
@@
where we expressed $\varphi_{n}=\sqrt{\pi/2E_{n}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right)$ and $\Pi_{n}=-\mathrm{i}\sqrt{E_{n}/2\pi}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ for $n\geq 1$ in bosonic operators $a^{\dagger}_{n}$, $a^{\vphantom{\dagger}}_{n}$ with $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$. Note that $\varphi_{0}$ (and zero modes in general) should be treated separately, but this will have no consequence here. In the SM [60], we also discuss SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8).
The time evolution of any observable, such as $\varphi(x)$, will contain terms like $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Thus, as shown in the SM [60], to exhibit PWT as defined by Eq. (1), all $u_{n}(x)$ must have definite parity, i.e., be invariant under reflections $x\mapsto-x$ up to a sign, which must equal the value of each phase factor $\mathrm{e}^{\pm\mathrm{i}E_{n}T}$ for a common $T>0$. Moreover, the $u_{n}(x)$ have definite parity if and only if $v(x)$, $K(x)$ [and $q(x)$ if non-zero] are even, i.e., $H_{\textrm{iTLL}}$ is reflection symmetric. In conclusion, PWT boils down to requirements on the spectrum:
An inhomogeneous TLL exhibits PWT if and only if $v(x)$ and $K(x)$ are even and all eigenvalues are of the form $E_{n}=\pi m_{n}/T$ for certain integers $m_{n}$.
From Properties (i) and (ii) and since $\lambda_{0}=0$ here, these integers $m_{n}$ satisfy $0=m_{0}<m_{1}<\ldots$ with $m_{n}$ even (odd) if $n$ is even (odd). Moreover, from Property (iv), they scale as $m_{n}\sim n$ for large $n$, i.e., $m_{n}=n+o(n)$.
Lastly, to probe PWT based on Eq. (5), we note that the results above allow one to compute correlation functions for inhomogeneous TLLs; see the SM [60]. In particular, the ground-state $\varphi\varphi$ correlation function is

### v3

@@ Abstract
The transfer of information from one part of a quantum system to another is fundamental to the understanding and design of quantum information processing devices. In the realm of discrete systems such as spin chains, inhomogeneous networks have been engineered that allow for the perfect transfer of qubits from one end to the other. Here, by contrast, we investigate the perfect transfer of information in continuous systems, phrased in terms of wave propagation. A remarkable difference is found between systems that possess conformal invariance and those that do not. Systems in the first class enjoy perfect wave transfer (PWT), explicitly shown for one-particle excitations and anticipated in general. In the second class, those that exhibit PWT are characterized as solutions to an inverse spectral problem. As a concrete example, we demonstrate how to formulate and solve this problem for a prototypical class of bosonic theories, showing the importance of conformal invariance for these theories to enjoy PWT. Using bosonization, our continuum results extend to theories with interactions, broadening the scope of perfect information transfer to more general quantum systems.
Introduction—A fundamental task in quantum information processing is to transfer information from one register to another [1, 2]. Optical fibers may be used, but efficiency can be improved using a quantum wire as both communication channel and registers [3]. Since this relies on collective phenomena to transfer quantum states, efficiency is optimized by engineering the system’s many-body properties. This allows not only for high-fidelity but even perfect transfer of states without dynamical control of individual sites. Such perfect state transfer (PST) was first shown theoretically for wires modeled as spin chains with inhomogeneous hopping [4, 5], then extended to spins on graphs and other networks [6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and demonstrated experimentally [16, 17, 18, 19, 20, 21]. However, PST has so far been addressed only in the discrete realm, while powerful continuum frameworks to describe collective quantum phenomena remain untapped.
@@
In this Letter, we harness that power to investigate quantum information transfer in inhomogeneous continuous systems in 1+1 dimensions (1+1D). At low energies, such examples are realizable using, e.g., spin ladders [22], Josephson-junction arrays [23], ultracold atoms [24], and nanowires [25, 26, 27]. In this framework, rephrasing using waves becomes natural, motivating the term perfect wave transfer (PWT): As for states in PST, a system on an interval [L/2,L/2][-L/2,L/2] of length $L$ with suitable boundary conditions (BCs) exhibits PWT if propagating any initial wave for some time $T$ amounts to a reflection; see Fig. 1.
Figure 1: (a) Illustration of PWT with Neumann BCs showing an initial wave that evolves to its reflection at a certain time $T$. (b) Its manifestation in an inhomogeneous CFT with a given profile $v(x)/v_{0}$ (inset), showing a localized initial wave propagating along curved light-cone trajectories that recombine at the reflected point at $t=T\equiv L/v_{0}$ given by Eq. (7). Specifically, the real part of boson two-point correlations (17) at the free-fermion point is plotted, which by bosonization encodes the same information as $\log F(t)$ in Eq. (5) for deltalike waves $\xi_{1}$ centered at 3L/8-3L/8 and $\xi_{2}$ at $x$ (setting $k_{\textrm{F}}=0$ for simplicity and $t^{\prime}=0$).
Concretely, there is PWT if
$\langle\mathcal{O}(x,0)\rangle=\langle\mathcal{O}(-x,T)\rangle$ (1)
for all time-evolved observables $\mathcal{O}(x,t)$ and expectation values $\langle\cdot\rangle$ with respect to any state. The dynamics will be time independent, but the problem is also interesting for driven systems [19, 20, 21], with randomness [28], or relaxed to a property for specific classes of observables or states. Ultimately, our motivation for understanding PWT is the efficient high-fidelity transfer of information—a problem that has attracted recent attention, e.g., in the design of distributed networks for quantum computation [29].
Related advances include nonequilibrium studies of inhomogeneous versions of conformal field theory (CFT), i.e., conformally invariant quantum field theories [30] deformed so that the usual constant propagation velocity becomes a profile $v(x)$ [31, 32, 33, 34, 35, 36]. Among their applications, such inhomogeneous CFTs effectively describe gapless quantum many-body systems with mesoscopic inhomogeneities, e.g., spin chains with spatially varying hopping and trapped ultracold atoms; cf. [37]. Special cases include the profile $v(x)=v\sqrt{1-(2x/L)^{2}}$ for some constant $v>0$ [38, 39]—the continuum counterpart of the inhomogeneous spin chain [4] first exhibiting PST—and profiles produced by Möbius transformations [40, 41, 42, 43, 44, 45, 46, 47]. Note that for PWT, as for other transport properties in finite systems, the choice of BCs [48, 49, 39, 50] is important.
Engineering PWT will, at heart, be a spectral problem, amounting to whether the propagator coincides with the parity operator at certain times. This is shown for prototypical quantum field theories describing an inhomogeneous quantum wire—our continuous channel—focusing on universal low-energy descriptions where conformal invariance can be parametrically turned on or off. Notably, these are directly applicable to ultracold atoms [51, 52, 49], while such experimental systems are even fundamentally continuous [53], in itself warranting a continuum treatment of information transfer. Remarkably, for bosonic theories [54, 55, 56, 57, 58, 59, 60, 61], we arrive at an inverse spectral problem for Sturm-Liouville operators. Its solution determines if the channel enjoys PWT, and we show that conformal invariance is important if the system is required to be sufficiently regular. Using bosonization [62, 63, 64, 65, 66, 67], our results extend PWT or PST to a large class of theories, including interacting fermions, and we also discuss a proposal to extend beyond low energies. Lastly, in addition to quantum communication in solid-state devices [68, 69], our PWT results relate to relativistic quantum information [70, 71], and potentially to communication over long distances [72].
Inhomogeneous CFT—To begin, we study the continuum limit of inhomogeneous spin chains [3, 4]. In the gapless regime, their continuous counterpart is an inhomogeneous CFT with a velocity profile $v(x)>0$ [31, 32, 33, 34, 35, 36]. The simplest example is that of inhomogeneous free fermions, given by the Hamiltonian (setting $\hbar=1$) [73]
$H_{\textrm{iFF}}\equiv\int_{-L/2}^{L/2}\hskip-2.3pt\mathrm{d}x\,\frac{v(x)}{2}\Bigl[\!\left.:\!\hskip-0.5pt\psi^{\dagger}_{+}(-\mathrm{i})\partial_{x}\psi^{\vphantom{\dagger}}_{+}\hskip-0.5pt\!:\right.+\left.:\!\hskip-0.5pt\psi^{\dagger}_{-}\mathrm{i}\partial_{x}\psi^{\vphantom{\dagger}}_{-}\hskip-0.5pt\!:\right.+\mathrm{h.c.}\Bigr]$ (2)
@@
PST was probed in [4] using a time-dependent overlap between one-particle states that were mutual reflections. Their continuum counterparts are states created by fields
@@
where $k_{\textrm{F}}$ denotes the Fermi momentum [79]. As a subset of possible states, consider
@@
for smooth enough functions $\xi_{j}(x)$, interpreted as “waves” of fermionic excitations in each state, and study the time-dependent overlap
@@
The states in Eq. (4) form a reduced class, as only one-particle excitations are possible, and right and left movers are equally distributed. (The latter restriction is lifted in the SM [74].)
To probe PWT for one-particle excitations as in [4], we restrict to specular waves, $\xi_{2}(x)=\xi_{1}(-x)$, and compare $F(t)$ in Eq. (5) with $\langle\Psi_{1}|\Psi_{1}\rangle$. Specifically, there is one-particle PWT if there exists $T>0$ so that
@@
for all initial waves $\xi_{1}(x)$ in Eq. (4) [80].
Eq. (6) amounts to a comparison between two-point correlation functions at equal and nonequal times. Such correlations can be computed in any inhomogeneous CFT by mapping them to the corresponding homogeneous correlations [36]. The details given in the SM [74] distill to comparing spacetime points after conformal transformations to coordinates
@@
The comparison’s outcome shows that reflection symmetry is essential: Inhomogeneous CFTs exhibit one-particle PWT if and only if $v(x)$ is even. Moreover, this transfer occurs at odd multiples $(2m+1)T$ (for $m=0,1,\ldots$) of the time $T=L/v_{0}$ for excitations to cross the system’s entire length. Note that $y\in[-L/2,L/2]$ if $v(x)$ is even.
@@
As an example, consider the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$, effectively describing the spin chain in [4]. It exhibits PWT, just like for any even profile $v(x)$, in contrast to [4], where that profile was exceptional in this respect. Notably, the square-root profile is positive everywhere except at $x=\pm L/2$. Thus, strictly speaking, $v(x)$ should be regularized before using the tools in [36], but since the integral defining $1/v_{0}$ in Eq. (7) converges, it has no consequence for PWT [81].
@@
Bosonic theories—The above result and observations should not come as a surprise. Indeed, if one compares discrete and continuous theories, then an inhomogeneous CFT can describe only low-energy excitations of a given lattice model. Moreover, we show [74] that any Hamiltonian (with pure point spectrum) must commute with the parity operator for PWT to be possible, i.e., $v(x)$ must be even, as should be expected. Furthermore, since excitations propagate ballistically in CFT, one might anticipate that this is not only necessary but also sufficient to ensure PWT with respect to all states. E.g., the model (2) has a linear spectrum and is free, suggesting that PWT for one-particle excitations extends to many particles. The rest of this Letter concerns to what extent this is true more generally: What are necessary and sufficient conditions for PWT as defined in Eq. (1)?
@@
To answer this question and understand the role of conformal invariance, we should not assume the latter. Specific nonconformal examples include 1+1D fermions with nonlocal interactions [82, 83, 84] or the Lieb-Liniger model [85]. While interesting in their own right, we address this question using a paradigmatic low-energy theory, namely, inhomogeneous Tomonaga-Luttinger liquids (TLLs) [54, 55, 56, 57, 58, 59, 60, 61]. By bosonization, this covers fermionic models generalizing the one in Eq. (2), while manipulating the inhomogeneities allows one to controllably turn conformal invariance on or off. The theory is formulated using a compactified bosonic field $\varphi(x)$ (modulo $2\pi$) with conjugate $\Pi(x)$ satisfying $\partial_{x}\varphi(x),\Pi(x^{\prime})]=\mathrm{i}\partial_{x}\delta(x-x^{\prime})$. The Hamiltonian is [73]
@@
$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\hskip-0.5pt\biggl(\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}\biggr)\hskip-0.5pt\!:\right$.\!, (8)
@@
where, in addition to $v(x)>0$, the Luttinger parameter $K(x)>0$ also depends on $x$. For simplicity, we assume that both functions are smooth. For $K(x)=K$ constant, Eq. (8) is an inhomogeneous CFT, and if $K=1$, it is the bosonized version of the free-fermion model (2) [62, 63, 64, 65, 66, 67]. However, $K^{\prime}(x)\neq 0$ locally couples right and left movers nontrivially; thus, they do not commute, and the theory is not conformal [61]. Again, we impose open BCs:
@@
$\hskip-2.84526pt\jmath(x)\big|_{x=\pm L/2}=0,\;\;\textnormal{where}\;\;\jmath(x)\equiv-\frac{v(x)K(x)}{\pi}\partial_{x}\varphi(x)$ (9)
@@
is the particle current. As a rule, we further assume that $K^{\prime}(x)$ is zero at $x=\pm L/2$ so that right and left movers are coupled only via the BCs (9) at those points [86].
@@
Sturm-Liouville (SL) theory—Following [87, 88, 89, 90, 91, 92, 37], we use SL theory to find eigenfunctions, using which the fields can be expanded and the Hamiltonian (8) diagonalized [93]. Indeed, using integration by parts and Eq. (9),
@@
$H_{\textrm{iTLL}}=\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}w(x)\!\left.:\!\hskip-0.5pt\Biggl(\biggl[\frac{\pi\Pi(x)}{w(x)}\biggr]^{2}+\varphi(x)\mathcal{A}\varphi(x)\Biggr)\hskip-0.5pt\!:\right$.\! (10)
@@
$w(x)\equiv\frac{K(x)}{v(x)},\quad p(x)\equiv v(x)K(x),\quad q(x)=0,$ (12)
where $\mathcal{A}$ acts on the space $\mathcal{D}$ of smooth [94] functions $u(x)$ on [L/2,L/2][-L/2,L/2] that satisfy $v(x)K(x)\partial_{x}u(x)\big|_{x=\pm L/2}=0$, i.e., Neumann BCs when $v(x),K(x)>0$.
@@
It is regular if certain conditions are satisfied [95], and irregular otherwise. This matters due to properties that hold for regular problems [96]:
@@
    The spectrum consists of eigenvalues $\lambda_{n}$ that are real and ordered so that $0\leq\lambda_{0}<\lambda_{1}<\ldots$.
@@
    Each $\lambda_{n}$ has a unique square-integrable eigenfunction $u_{n}(x)$ with $n$ isolated zeros in [L/2,L/2][-L/2,L/2].
@@
    The $u_{n}(x)$ form a complete orthonormal basis for square-integrable functions on [L/2,L/2][-L/2,L/2] with weight $w(x)$, i.e., $\int_{-L/2}^{L/2}\mathrm{d}x\,w(x)u_{n}(x)u_{n^{\prime}}(x)=\delta_{n,n^{\prime}}$.
Furthermore, assuming that the integral defining $1/v_{0}$ in Eq. (7) converges, Weyl’s law implies the following:
@@
    $\lambda_{n}$ grows at most as $n^{2}$ for $n\gg 1$, i.e., $\lambda_{n}=O(n^{2})$;
cf. the SM [74] and [97, 98].
@@
$\varphi(x)=\sum_{n=0}^{\infty}\varphi_{n}u_{n}(x),\qquad\Pi(x)=\sum_{n=0}^{\infty}\Pi_{n}w(x)u_{n}(x)$ (14)
with $\bigl[\varphi_{n^{\vphantom{\prime}}},\Pi_{n^{\prime}}\bigr]=\mathrm{i}\delta_{n,n^{\prime}}$ and thereby diagonalize $H_{\textrm{iTLL}}$ in Eq. (8). Indeed, the latter becomes
@@
where we expressed $\varphi_{n}=\sqrt{\pi/2E_{n}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right)$ and $\Pi_{n}=-\mathrm{i}\sqrt{E_{n}/2\pi}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ for $n\geq 1$ in bosonic operators $a^{\dagger}_{n}$, $a^{\vphantom{\dagger}}_{n}$ satisfying $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$ [99].
[…]
The SM [74] also examines operators of the form (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8).
The time evolution of any observable, such as $\varphi(x)$, contains terms like $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Thus, for Eq. (1) to hold for all states, each $u_{n}(x)$ must have definite parity, i.e., be invariant under reflections $x\mapsto-x$ up to a sign, which must equal each phase factor $\mathrm{e}^{\pm\mathrm{i}E_{n}T}$ for a common shortest possible $T>0$. Moreover, the $u_{n}(x)$ have definite parity if and only if $v(x)$, $K(x)$ [and $q(x)$ if nonzero] are even, i.e., $H_{\textrm{iTLL}}$ is reflection symmetric. This leads to a statement in terms of spectral requirements [74]: Inhomogeneous TLLs exhibit PWT as defined in Eq. (1) if and only if $v(x)$ and $K(x)$ are even and the eigenenergies are $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ so that [100]
  1. (a)
    $0=m_{0}<m_{1}<\ldots$ with each $m_{n}$ even or odd if $n$ is even or odd,
  2. (b)
    $m_{n}\sim n$ for large $n$, i.e., $m_{n}=n+o(n)$.
To probe PWT based on Eq. (5), note that the results above allow one to compute correlations for inhomogeneous TLLs [74]. In particular, the ground-state $\varphi\varphi$ correlation function is [99]